OpenStudy (anonymous):

mark, jessica & nate each downloaded music from the same website. mark downloaded 10 songs in total consisting of pop, rock, & hip hop. jessica downloaded 5 times as many pop songs, twice as many rock and 3 times as many hip hop songs as mark. she downloaded 28 songs total. if nate downloaded 20 songs total with 3 times as many pop songs, 3 times as many rock songs, and the same number of hip hop songs as mark, how many songs of each type did mark download? the answer is Mark downloaded 1 pop songs, 4 rock songs and 5 hip hop songs, correct?

OpenStudy (anonymous):

Can you tell me how you got that?

OpenStudy (anonymous):

2(p + r + h = 10) -(5p + 2r + 3h = 28) +(3p + 3r + h = 20) 3r = 12 r = 4

OpenStudy (anonymous):

Good job :)

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

It is right, if you want to go back and check it you can replace p,r, and h with the values you found and make sure you get 10, 28, and 20

OpenStudy (anonymous):

i did. c:

OpenStudy (anonymous):

Good job :)

OpenStudy (anonymous):

thank you! lol.